The velocity of waves on a string is 95.0 m/s. If the frequency of standing waves is 160 Hz, how far apart are two adjacent nodes? If two successive overtones of this vibrating string are 240 Hz and 320 Hz, what are the frequency of the fundamental and the length of the string?
Answer #1

velocity v = 95 m/s

(a)

frequency n = 160 Hz

Wavelength λ = v/n

                     = 95/160 =  0.593 m

Distance between two adjacent nodes d = λ/2 = 0.296 m

(b)

Fundamental frequency n = 320 - 240 = 80 Hz

Length of the string l = (1/2n) v = 0.593 m